THE FOLLOWING IS THE DEMAND FUNCTION FOR SUGAR:

Qs = 20 -5Ps + 3Pe + 6Y

Where             Qs                    = demand for sugar (in pounds)

            Ps        = $5     = price of “sugar” for pounds

            Pe         = $100 = price of “equal” pounds

                        Y        = $200 = per capital income for a week

  1. What is the demand for sugar Qs?
  2. Compute and interpret (using a one percent change) the price elasticity.
  3. Compute and interpret (using a one percent change) the cross price elasticity.

Compute and interpret (using a one percent change) the income elasticity.
Answer #1

Demand : Qs = 20 - 5 x 5 + 3 x 100 + 6 x 200 = 1495

Price elasticity = dQs/Ps x Ps/Qs = -5 x 5/1495 = - 0.017

Cross price elasticity = dQs/dPe x Pe/Qs = 3 x 100/1495 = 0.2

Income elasticity = dQs/dY x Y/Qs = 6 x 200/1495 = 0.8